The legs of a right triangle are 33ft & 56ft. Need to compute the length of the hypotenuse.
Answer by algebrahouse.com(1659)

Pythagorean Theorem is a² + b² = c² {a and b are the legs, c is the hypotenuse}

33² + 56² = c² {substituted into the pythagorean theorem}
1,089 + 3,136 = c² {squared the 33 and 56}
4,225 = c² {added}
c = 65 {took square root of both sides}

hypotenuse = 65 ft

For more help from me, visit: www.algebrahouse.com